OpenStudy (anonymous):

how would I find the derivative of log(base10)(t+2^t) with respect to t? not sure how I would go about a differentiation problem like this.

OpenStudy (anonymous):

\[[\log_{10} (t+2^{t})] dt\] ?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

\[[1+\ln 2\times(2^{t}) ]/[\ln10\times (t+2^{t})]\]

OpenStudy (anonymous):

what differentiation rules did you use? if you don't mind me asking

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Chain_rule